Show that βˆ‡π‘Ÿn = π‘›π‘Ÿn-2"\vec r"


1
Expert's answer
2021-08-19T15:27:17-0400

"\triangledown r^{n}=nr^{n-2}\vec r"

"\vec r=x\hat i+y\hat j+z\hat k"

"|\vec r|^{n}=r^{n}=\sqrt [n] {x^{2}+y^{2}+z^{2}}"

"\triangledown=\frac{d}{dx}\hat i+\frac{d}{dy}\hat j+\frac{d}{dz}\hat k"

"\triangledown r^{n}=\frac{dr^{n}}{dx}\hat i+\frac{dr^{n}}{dy}\hat j+\frac{dr^{n}}{dz}\hat k"

"\frac{dr^{n}}{dx}=\frac{d(x^{2}+y^{2}+z^{2})^{\frac{n}{2}}}{dx}"

"=\frac{n}{2}(x^{2}+y^{2}+z^{2})^{{\frac{n}{2}}-1}\frac{d}{dx}(x^{2}+y^{2}+z^{2})"

"=\frac{n}{2}(x^{2}+y^{2}+z^{2})^{{\frac{n}{2}}-1}2x"

"=n(x^{2}+y^{2}+z^{2})^{{\frac{n}{2}-1}}x"

"||y"

"\frac{dr^{n}}{dy}=\frac{d(x^{2}+y^{2}+z^{2})^{\frac{n}{2}}}{dy}"

"=n(x^{2}+y^{2}+z^{2})^{\frac{n}{2}-1}y"

And;

"\frac{dr^{n}}{dz}=\frac{d(x^{2}+y^{2}+z^{2})^{\frac{n}{2}}}{dz}"

"=n(x^{2}+y^{2}+z^{2})^{\frac{n}{2}-1}z"

Hence;

"\triangledown r^{n}=n(x^{2}+y^{2}+z^{2})^{\frac{n}{2}-1}(x\hat i+y\hat j+z\hat k)\newline \triangledown r^{n}=n(x^{2}+y^{2}+z^{2})^{\frac{n}{2}-1}\vec r"

"\triangledown r^{n}=n(x^{2}+y^{2}+z^{2})^{\frac{n-2}{2}}\vec r"

Thus,

"\triangledown r^{n}=nr^{n-2}\vec r"